OpenStudy (anonymous):

Find the derivative: y=arcsin(x/3)

OpenStudy (rogue):

Apply the chain rule. The derivative of arcsine is\[\frac {d}{dx} \sin^{-1} (u) = \frac {u'}{\sqrt{1-u^2}}\]

OpenStudy (anonymous):

Is it \[1/\sqrt{9-x^2}\]?

OpenStudy (rogue):

Yeah, nice job.

OpenStudy (anonymous):

thanks for your help!